15-E1. Demand Q is negatively associated with price P. for every increase of $5 in price, demand decreases by 40 units. When price is $10, demand demand is 50 units. What is the demand equation?
Answer #1

Slope of demand equation = change in quantity demanded / change in price = -40/5 = -8

Thus demand equation is

Qd = -8P + a

Where Qd is quantity demanded, P is price and a is intercept. Also given, when price is 10, demand is 50. Substituting in equation, we have

50 = -8(10) + a

a = 130

Thus demand equation is

Qd = -8P + 130